OpenStudy (anonymous):

Determine the equations of both lines that are tangent to the graph of f(x) = x^2 and pass through the point (1,-3).

myininaya (myininaya):

f'(x)=2x f'(a)=2a so the equation of the line is y=2ax+b we know a point on this line (1,-3) so we -3=2a(1)+b -3=2a+b ....

myininaya (myininaya):

thinking...

OpenStudy (anonymous):

One would have a negative slope, the other would be positive

myininaya (myininaya):

right

OpenStudy (anonymous):

Would those two lines be orthogonal to each other?

myininaya (myininaya):

that looks possible it looks like it could form a 90 degree angles

OpenStudy (anonymous):

Ok so then they must have reciprocal slopes

myininaya (myininaya):

opposite reciprocal slopes

myininaya (myininaya):

wait i htink i'm fixing to figure something out....

myininaya (myininaya):

now we need to figure the y intercept for each line which is easy since we know a point on both lines. both lines have the point (1,-3)

myininaya (myininaya):

y=6x+b (1,-3) -3=6(1)+b -3=6+b -3-6=b -9=b so y=6x-9 is the line that is tangent to the point (3,9) on the curver y=x^2 that passes through (1,-3) y=-2x+b (1,-3) -3=-2(1)+b -3+2=b -1=b so y=-2x-1 is the line tangent to the point (-1,1) on the curve y=x^2 that passes through (1,-3)

myininaya (myininaya):

we can check if you want Let's find the tangent line at (3,9) and see if passes through (1,-3) so f'(x)=2x f'(3)=6 y=6x+b 9=6(3)+b 9-18=b b=-9 so y=6x-9 is (1,-3) on that line -3=6(1)-9=-3 so yes! now let's check the tangent line at (-1,1) and see if passes through (1,-3) so f'(x)=2x f'(-1)=-2 y=-2x+b 1=2+b b=-1 y=-2x+-1 is (1,-3) on that line -3=-2(1)+-1=-3 so yes! :) we are finished!

myininaya (myininaya):

so the lines arent orthogonal

myininaya (myininaya):

do you have any questions?

OpenStudy (anonymous):

One question

OpenStudy (anonymous):

Why did you choose the general points (a,a^2) to be X1,Y1 respectively? Does it make a difference if I were to use them as X2,Y2 in the slope formula?

myininaya (myininaya):

(a^2-(-3))/(a-1)=(-3-a^2)/(1-a) we can make both sides look the same (-1)/(-1)=1 so if we multiply our second fraction by (-1)/(-1) we would get the first fraction or vice versa

myininaya (myininaya):

(y1-y2)/(x1-x2)=(y2-y1)/(x2-x1) (-1)/(-1)*(y1-y2)/(x1-x2)=(y2-y1)/(x2-x1)

myininaya (myininaya):

a-b is the same as -(b-a)

myininaya (myininaya):

(6-2)/(3-1)=4/2=2 (2-6)/(1-3)=-4/(-2)=2

myininaya (myininaya):

cool?

OpenStudy (anonymous):

Yes thank you very much. I have just a few more questions if you don't mind helping me out with them.

myininaya (myininaya):

go for it

OpenStudy (anonymous):

If f(a) = 0 and f'(a) = 6, find the limit h --> 0 of: f(a + h) ------- 2h

myininaya (myininaya):

i got it scanning it now

myininaya (myininaya):

i just used the definition of derivative

myininaya (myininaya):

and plugged and what they gave me and solved for limh->0 (f(a+h)/(2a))

myininaya (myininaya):

do you understand both problems we did?

myininaya (myininaya):

oops solved for limh->0(f(a+h)/(2h))

OpenStudy (anonymous):

I don't understand how you did the second question.

myininaya (myininaya):

what is definition of the derivative at a of f

OpenStudy (anonymous):

Wait, ok wouldn't it be 12 for the answer because: 6 = f(a+h)/h * (1/2) 6(2) = f(a+h)/h

OpenStudy (anonymous):

Sorry I don't know what happened

myininaya (myininaya):

doesn't the question above ask for lim h->0 (f(a+h)/(2a))

myininaya (myininaya):

oops the bottom is suppose to read 2h

myininaya (myininaya):

so we have 6=limh->0(f(a+h)/(2h))

myininaya (myininaya):

oops 6=limh->0(f(a+h)/h) do you understand this far?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok but the want 1/2* limh->0(f(a+h)/h) so if we multiply the above equation on one side by 1/2 don't we have to do it to the other side?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok 6*(1/2)=3

OpenStudy (anonymous):

Got it. Thank you.

OpenStudy (anonymous):

Ok last question for now is this: Determine the value of "a', given that the line "ax - 4y + 21 = 0" is tangent to the graph y = a/x^2 at x = -2

myininaya (myininaya):

if you write that tangent line in y=mx+b form we can find the slope of the line

myininaya (myininaya):

so the equation of that line can be written y=ax/4-21/4 where a/4 is the slope of this tangent line

OpenStudy (anonymous):

The slope of the line is a right?

OpenStudy (anonymous):

ok so then we equate the two equations together

OpenStudy (anonymous):

right?

myininaya (myininaya):

so you are talking about finding the derivative of that function y=a/x^2 at x=-2 and then setting that equal to the slope of tangent line and solving for a? that's what I'm working on right now

OpenStudy (anonymous):

yes and I got 7 when equating the two equations together. But I don't understand how we can equate a y with a y' (derivative)

myininaya (myininaya):

so we get a=a so this equation holds for any a is what I get

OpenStudy (anonymous):

Oh so you are making them equal to each other by solving for a?

OpenStudy (anonymous):

a = , a =, and then equating them?

myininaya (myininaya):

derivative means slope this slope is suppose to be the same as the slope tangent to whatever point we are talkng about so we can find the slope of the tangent line and find the derivative of the curve at the given point and set them equal to find what a should be but the resulting equation is a=a this equation holds for every number 3=3 4=4 it is never not true

myininaya (myininaya):

a/4=-2a/(x^3) at x=-2 is what I did

OpenStudy (anonymous):

how did you get a/4?

myininaya (myininaya):

a/4 is the slope of the tangent line

OpenStudy (anonymous):

OOHHHH I understand So we can take a/4 to be the slope, which must be the same as the graph's slope at that point?

myininaya (myininaya):

right just like the very very question we did sort of

OpenStudy (anonymous):

Got it. Thank you so much for your help. I have to go now. I really appreciate you taking your time to do this for me.

myininaya (myininaya):

no problem become my fan :)

OpenStudy (anonymous):

Sure, see you.

myininaya (myininaya):

later